OpenStudy (anonymous):

Without using calculator, can somebody tell how to solve √34700, using this information: √3.47=1.863 and √34.7=5.891. Please reply quickly!

OpenStudy (cwrw238):

√34700 = √3.47 * √10,000 = 100*√3.47 = 186.3

OpenStudy (anonymous):

Thank you so much! :)